OpenStudy (anonymous):

I am needing help! Find the volume of a box that measures 4 1/3 feet by 1 1/2 feet by 3 1/3 feet

OpenStudy (anonymous):

volume = a x b x c

OpenStudy (jdoe0001):

volume of a cube = length * width * height thus volume =\(\bf 4\frac{1}{2}\times 1\frac{1}{2}\times 3\frac{1}{3}\)

OpenStudy (anonymous):

where a,b,c are edge lenths

OpenStudy (anonymous):

13 x 3 x 10 = 490

OpenStudy (anonymous):

would the denominator be 6

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

13x 3 x 10 = 390?